OpenStudy (anonymous):

if a divide w and w divide k, then a divde k (z and y is intergers) proof plz I got up to az = w and wy = k a(zy) = k and I am stucked here

OpenStudy (aum):

\[ \frac wa = z; ~~ a = \frac wz \\ \frac kw = y; ~~ k = wy \\ \frac ka = \frac {wy}{w/z} = wy * \frac zw = zy ~~~~~\text {zy is integer because z and y are integers.} \]

OpenStudy (anonymous):

thx

OpenStudy (aum):

yw.